# Discrete Max-Focusing

### v2

@@ 1 Introduction
4. The QFC, and associated statements like the Bousso bound and the GSL, exhibit some apparent redundancies. For example, they can equally be applied to the interior and to the exterior of a cut of the horizon of a black hole; their von Neumann expansions are obviously equal and opposite. When the von Neumann entropy is replaced by conditional min and max entropies, these redundancies are not resolved: the future-outward max expansion of the interior is minus the past-outward min expansion of the exterior; and the future-outward min expansion of the interior is minus the past-outward max expansion of the exterior; and similarly for the future-inward/past-outward pairings. This leads to ambiguities: the GSL was introduced as a statement about the generalized entropy of the black hole exterior, but apparently applies equally well to the interior.
Figure 1: Generalized entanglement wedge $e(a)$ of a “boosted crescent” wedge $a$ in 2+1-dimensional Minkowski space. The tips of the crescent are null separated, and the edge of $e(a)$ contains the null geodesic that connects these two points. On this null portion, a numerical expansion is not defined and hence cannot be said to vanish. Thus, it is not correct to say that $e(a)$ is “extremal” (in the sense of vanishing expansion) on the edge portion ð$\eth e(a)\setminus\eth a$. However, our discrete notion of nonexpansion, Def. 28, does hold on this edge portion. Moreover, the notion of “extremal” can be replaced with the more precise and robust notion of “throat,” which also holds here; see Def. 50 and Theorem 61.
@@
Thus our reformulation minimizes the axiomatic foundation of the holographic properties of semiclassical spacetime, while maximizing its reach. In particular, the fourth shortcoming noted above — the redundancy inherent in the traditional formulation of the QFC — is also addressed. An inward max-expansion is no longer defined, nor is an outward min-expansion defined. Only the outward max-expansions are explicitly defined, and not as a numerical quantity, but only through the qualitative notion of outward max-nonexpansion. (Inward min-expansion remains implicitly defined through the outward max-expansions of the complement wedge, but for conceptual clarity only the latter is referred to in this paper. The definition of the generalized min entanglement wedge as the complement of a generalized max entanglement wedge of a suitable complement region will be further developed elsewhere.)
Our work has two strands that may seem unrelated. The first strand is the generalization of the classical notion of nonexpansion to nonsmooth surfaces such as those resulting from intersections and caustics of null congruences. The second strand is the reduction of the types of entropies that appear in definitions to the max conditional entropy only. We would present these results in separate manuscripts, were it not for the fact that the two strands come together in the final result that only outward deformations of wedges should be considered. Classical nonexpansion is about what happens next as we enlarge a wedge; our definition of “nonexpanding” cannot be reproduced by considering inward deformations instead. Meanwhile, the outward-directedness of the conditional max entropy is inherent both in its mathematical properties — the max and min entropies obey opposite chain rule inequalities that work well with unions and intersections, respectively — and in its physical interpretation as measuring the resources needed to reconstruct semiclassical data in a larger region from a more fundamental algebra of a smaller one Bousso:2022hlz ; Bousso:2023sya .
@@
In Sec. 3.1, we define discrete notions of future-nonexpanding and past-nonexpanding, in terms of the negative sign of the generalized smooth max entropy of outward deformations of a wedge $a$, conditioned on the original $a$. Importantly, these definitions make sense at all points on ð$\eth a$, whether they lie on null portions, corners, or smooth spacelike portions of ð$\eth a$. We formulate a discrete, max-version of the Quantum Focusing Conjecture based on these definitions. In Sec. 3.2 we establish some immediate implications of the definitions and conjecture: the fact that null portions of an edge are past- and future-nonexpanding; the persistence of nonexpansion along lightsheets; the Bousso bound; and the Generalized Second Law of thermodynamics. In Sec. 3.4, we identify circumstances under which the expansion can be assigned a continuous value $\Theta^{\pm}$, and we prove relations between the discrete and continuous notions of nonexpansion.
@@
In Sec. 5, we re-develop important concepts and results in semiclassical gravity and holography, using only the criterion of Discrete Max-Nonexpansion instead of a numerical quantum expansion, and using Discrete Max-Focusing in place of the original QFC. In Sec. 5.1 we introduce the concepts of an antinormal wedge (one that is both past- and future-nonexpanding), and of a normal wedge (one whose complement is antinormal). We prove that the union of antinormal wedges is antinormal.
In Sec. 5.2 we define marginal wedges. Traditionally these correspond to wedges in which at least one quantum expansion vanishes, so that the edge of the wedge lies on an apparent horizon. If both of the quantum expansions vanish, the edge is traditionally called a quantum extremal surface. In our discrete setting, we do not wish to start by imposing numerical conditions like $\Theta^{+}=0$, since a numerical expansion is not always defined. Nevertheless, an appropriate notion of marginal wedge can be recovered: namely, as an antinormal wedge that admits no future-outward deformation (or no past-outward deformation) that is antinormal. In fact this property implies something more specific and more relevant than just the vanishing of expansions: it defines the notions of stably future- and past-marginal Engelhardt:2018kcs , and of throat (as opposed to bulge or bounce) wedges Brown:2019rox ; Engelhardt:2023bpv . We prove that at points where a stably future-marginal wedge has a numerical value of the past max-expansion, it vanishes. Similarly, both numerical expansions of throat wedges vanish where they exist.
@@ 2.1 Wedges and Causal Structure
Let $s\subset M$ [^3] In this paper, $\subsetneq$ denotes a proper subset; $\subset$ permits equality. We use $\operatorname{int}s$, $\operatorname{cl}s$, and $\partial s$ to denote the interior, the closure, and the boundary of $s$ in $M$.
@@
The edge ð$\eth a$ and Cauchy horizons $H^{\pm}(a)$ of a wedge $a$ are defined by
ð$\eth a$ $\equiv\partial a\setminus I(a)~{},$ (3)
@@
ð$\eth a$ is closed in $M$ since ð$\eth a=M\setminus I(a)\setminus I(a^{\prime})$. Moreover, ð$\eth a$ is an immersed $C^{0}$ submanifold of codimension 2. However, ð$\eth a$ need not be embedded (for example, in $2$ spatial dimensions, suppose that $a$ is the union of two disjoint wedges whose edges intersect at a point).
@@
The area of ð$\eth a$ will be denoted $\operatorname{Area}(a)$.
@@
Given a wedge $a$, we distinguish between its edge ð$\eth a$ in $M$ and its edge $\delta\bar{}\hskip 1.99997pta$ as a subset of the conformal completion $\tilde{M}$ Wald:1984rg . If $a$ is asymptotic, the latter set may contain an additional piece, the conformal edge
ð~$\tilde{\eth}a\equiv\delta\bar{}\hskip 1.99997pta\cap\partial\tilde{M}~{}.$ (8)
@@ 3.1 Definitions
We emphasize that no numerical value is assigned to the expansion by the present definition; see Def. 39 below. To define past-nonexpanding, $H^{-}$ replaces $H^{+}$ in Eq. (21).
Figure 2: Illustration of Definition 28: a wedge $a$ is future-nonexpanding at $p$ if there exists an open neighborhood $O$ of $p$ such that all future-outward deformations of ð$\eth a$ within $O$ (green) do not increase the area (classically), or (semiclassically) have nonpositive smooth max-generalized entropy conditioned on $a$.
Figure 3: Discrete Max-Focusing. If the wedge $a$ is future-nonexpanding on a portion ð$\eth a^{+}$ of its wedge ð$\eth a$, then $a$ can be outward deformed along the future lightsheet emanating from ð$\eth a^{+}$ into wedges $b$, $c$ that obey $H_{\rm max,gen}(c|b)\leq 0~{}.$
@@
Let ð$\eth a^{+}$ (ð$\eth a^{-}$) denote the set of points $p\in\eth a$ where the wedge $a$ is future- (past-) nonexpanding.
@@
ð$\eth a^{+}$ and ð$\eth a^{-}$ are open subsets of ð$\eth a$ in the induced topology of ð$\eth a$.
@@
Let $p\in\eth a^{+}$. We must show that ð$\eth a^{+}$ contains an open subset that contains $p$. Let $O$ be the open set guaranteed to exist by Def. 28, and let $q\in\eth a\cap O$. Then $O$ contains $q$ and satisfies Eq. (21). Hence all points in ð$\eth a\cap O$ are future-nonexpanding: ð$\eth a\cap O\subset\eth a^{+}$. ∎
@@
Let $a$ be a wedge and let ð$\eth a^{+}$ be the set of points where $a$ is future-nonexpanding. The future-lightsheet of $a$ is the null hypersurface
@@
Let $a$, $b$, and $c$ be wedges such that $a\subset b\subset c$, and suppose that ð$\eth b\cup\eth c\subset\eth a\cup L^{+}(a)$. Then
@@
The same statement holds in the past direction, i.e., if ð$\eth b\cup\eth c\subset\eth a\cup L^{-}(a)$.
@@ 3.2 Immediate Implications
By assumption there exists an open neighborhood $O$ of $p$ in which the past and future null vectors normal to ð$\eth a$ are tangent to ð$\eth a$. None of the null generators of $H^{\pm}(a^{\prime})$ start on this portion. Since $a$ cannot be deformed along $H^{\pm}(a^{\prime})$ at all, there exists no wedge $b$ for which Eq. (21) could fail. ∎
Figure 4: Applications of Definition 28. All wedges shown are understood to be subsets of Minkowski spacetime. Left: the wedge $a$ is nonexpanding at all the dark blue points, including at the (concave) corner (blue dot). Center: $a$ is future-nonexpanding except at the (convex) corner, where the green deformation ð$\eth a\to\eth b$ fails to satisfy Eq. (21). Right: the time direction is suppressed, but all three spatial directions are now displayed. The edge ð$\eth a$ is now shown as a shaded surface. The corner (red dot) is neither convex nor concave. Such a corner fails to be future-nonexpanding, because in any open neighborhood (orange), the green deformation increases the area.
@@
Let $a$ and $b$ be wedges such that $a\subset b$, and suppose that ð$\eth b\subset\eth a\cup L^{+}(a)$. Assuming Discrete Max-Focusing, Conjecture 32,
ð$\eth b\cap L^{+}(a)\subset\eth b^{+}~{}.$ (24)
The same statement holds in the past direction: $+$ can be replaced by -- throughout.
@@
By Lemma 30, ð$\eth a\setminus\eth a^{+}$ is closed in the induced topology of ð$\eth a$. Hence ð$\eth b\cap L^{+}(a)=\eth b\setminus(\eth a\setminus\eth a^{+})$ is open in the induced topology of ð$\eth b$. Therefore, there exists a set $O$ that is open in $M$ such that ð$\eth b\cap O=\eth b\cap L^{+}(a)$.
Let $c\supset b$ such that ð$\eth c\subset\eth b\cup[H^{+}(b^{\prime})\cap O]$. Since $H^{+}(b^{\prime})\subset H^{+}(a^{\prime})$, we have ð$\eth c\subset\eth a\cup L^{+}(a)$ and hence the assumptions of Discrete Max-Focusing, Conjecture 32, are satisfied. The Conjecture then implies Eq. (25). ∎
@@
Let $a\subset c$ such that ð$\eth c\subset\eth a\cup L^{+}(a)$. That is, $c$ is a future-outward deformation of $a$ along orthogonal lightrays originating on the future-nonexpanding portion of its edge ð$\eth a$. Then
@@
The same statement also holds if ð$\eth c\subset\eth a\cup L^{-}(a)$.
@@
Let ${\cal C}$ be the boundary of the past of a subset $s$ of the conformal boundary of an arbitrary spacetime $M$. (Note that $s$ can be, but need not be, a single point.) If there exists a Cauchy slice $\Sigma$ of $M$ such that any wedge $a$ with ð$\eth a\subset{\cal C}\cap J^{+}(\Sigma)$ is past-nonexpanding, then $\cal C$ is called a future-causal horizon.
@@
Let $\cal C$ be a future-causal horizon. Let $b\subset c$ be wedges whose edges both lie on $\cal C$ with ð$\eth c\subset L^{-}(b)$. Then

### v3

@@ 1 Introduction
Figure 1: Generalized entanglement wedge $e(a)$ of a “boosted crescent” wedge $a$ in 2+1-dimensional Minkowski space. The tips of the crescent are null separated, and the edge of $e(a)$ contains the null geodesic that connects these two points. On this null portion, a numerical expansion is not defined and hence cannot be said to vanish. Thus, it is not correct to say that $e(a)$ is “extremal” (in the sense of vanishing expansion) on the edge portion $\eth e(a)\setminus\eth a$. However, our discrete notion of nonexpansion, Def. 28, does hold on this edge portion. Moreover, the notion of “extremal” can be replaced with the more precise and robust notion of “throat,” which also holds here; see Def. 51 and Theorem 62.
@@
Thus our reformulation minimizes the axiomatic foundation of the holographic properties of semiclassical spacetime, while maximizing its reach. An inward max-expansion is no longer defined, nor is an outward min-expansion defined. Only the outward max-expansions are explicitly defined, and not as a numerical quantity, but only through the qualitative notion of outward max-nonexpansion. (Inward min-expansion remains implicitly defined through the outward max-expansions of the complement wedge, but for conceptual clarity only the latter is referred to in this paper. The definition of the generalized min entanglement wedge as the complement of a generalized max entanglement wedge of a suitable complement region will be further developed elsewhere BoussoKaya .)
Our work has two strands that may seem unrelated at first. The first strand is the generalization of the classical notion of nonexpansion to nonsmooth surfaces such as those resulting from intersections and caustics of null congruences. The second strand is the reduction of the types of entropies that appear in definitions to the max conditional entropy only. We would present these results in separate manuscripts, were it not for the fact that the two strands come together in the final result that only outward deformations of wedges should be considered. Classical nonexpansion is about what happens next as we enlarge a wedge; our definition of “nonexpanding” cannot be reproduced by considering inward deformations instead. Similarly, unlike the von Neumann entropy, the conditional max entropy is also intrinsically outward-directed. This is inherent in its mathematical properties: the max and min entropies obey opposite chain rule inequalities that work well with unions and intersections, respectively. It is also inherent in its physical interpretation as measuring the resources needed to reconstruct semiclassical data in a larger region from a more fundamental algebra of a smaller one Bousso:2022hlz ; Bousso:2023sya . It is interesting that these seemingly unrelated considerations both give preference to outward, rather than inward, deformations of wedges.
@@
In Sec. 3.1, we define discrete notions of future-nonexpanding and past-nonexpanding, in terms of the negative sign of the generalized smooth max entropy of outward deformations of a wedge $a$, conditioned on the original $a$. Importantly, these definitions make sense at all points on $\eth a$, whether they lie on null portions, corners, or smooth spacelike portions of $\eth a$. We formulate a discrete, max-version of the Quantum Focusing Conjecture based on these definitions. In Sec. 3.2 we establish some immediate implications of the definitions and conjecture: the fact that null portions of an edge are past- and future-nonexpanding; the persistence of nonexpansion along lightsheets; the Bousso bound; and the Generalized Second Law of thermodynamics. In Sec. 3.4, we identify circumstances under which the expansion can be assigned a continuous value $\Theta^{\pm}$, and we prove relations between the discrete and continuous notions of nonexpansion.
@@
Therefore Discrete Max-Focusing implies, in the respective limits, all further consequences of the conditions listed in the previous two paragraphs. Penrose’s singularity theorem Penrose:1964wq follows from the present work in the regime in which it is formulated (the classical limit, in which the Null Energy Condition emerges from Discrete Max-Focusing). Similarly, Wall’s quantum singularity theorem Wall:2010jtc follows in its own regime, the limit where numerical von Neumann quantum expansions exist and agree (at least in sign) with the max-expansions. Other such statements include the area theorem Hawking:1971vc , the Gao-Wald theorem Gao:2000ga , and causal wedge inclusion for von-Neumann entanglement wedges Wall:2012uf .
An important future goal will be to formulate all important classical and semiclassical gravity theorems directly in the fully general language of Discrete Max-Focusing (rather than recovering them as limits), and to adapt their proofs. In the present work, we initiate this effort. We define and we prove vital properties of wedges and entanglement wedges, as fully general theorems in the language of Discrete Max-Focusing. Elsewhere, a robust singularity theorem Bousso:2025xyc has already been formulated and proven at the full level of generality and minimality of ingredients advocated here. The Discrete Max formulation and proof of other theorems, such as Gao-Wald or causal wedge inclusion, will be pursued in separate works.
In Sec. 5, we re-develop some important concepts and results in semiclassical gravity and holography, using only the criterion of Discrete Max-Nonexpansion instead of a numerical quantum expansion, and using Discrete Max-Focusing in place of the original QFC. In Sec. 5.1 we introduce the concepts of an antinormal wedge (one that is both past- and future-nonexpanding), and of a normal wedge (one whose complement is antinormal). We prove that the union of antinormal wedges is antinormal.
In Sec. 5.2 we define marginal wedges. Traditionally these correspond to wedges in which at least one quantum expansion vanishes, so that the edge of the wedge lies on an apparent horizon. If both of the quantum expansions vanish, the edge is traditionally called a quantum extremal surface. In our discrete setting, we do not wish to start by imposing numerical conditions like $\Theta^{+}=0$, since a numerical expansion is not always defined. Nevertheless, an appropriate notion of marginal wedge can be recovered: namely, as an antinormal wedge that admits no future-outward deformation (or no past-outward deformation) that is antinormal. In fact this property implies something more specific and more relevant than just the vanishing of expansions: it defines the notions of stably future- and past-marginal Engelhardt:2018kcs , and of throat (as opposed to bulge or bounce) wedges Brown:2019rox ; Engelhardt:2023bpv . We prove that at points where a stably future-marginal wedge has a numerical value of the past max-expansion, it vanishes. Similarly, both numerical expansions of throat wedges vanish where they exist. (Since quantum extremal surfaces obtained by Wall’s maximin construction Wall:2012uf are throats Brown:2019rox , this result implies that the standard QES prescription becomes a good approximation to the max- and min- entanglement wedges we define, in those settings where the von Neumann entropy provides a good approximation. This includes in particular the derivation of the Page curve for evaporating black holes from entanglement islands Penington:2019npb ; Almheiri:2019hni ).
@@ 2.1 Wedges and Causal Structure
Let $s\subset M$. [^3] In this paper, $\subsetneq$ denotes a proper subset; $\subset$ permits equality. We use $\operatorname{int}s$, $\operatorname{cl}s$, and $\partial s$ to denote the interior, the closure, and the boundary of $s$ in $M$.
@@
The edge $\eth a$ and Cauchy horizons $H^{\pm}(a)$ of a wedge $a$ are defined by
$\eth a$ $\equiv\partial a\setminus I(a)~,$ (3)
@@
$\eth a$ is closed in $M$ since $\eth a=M\setminus I(a)\setminus I(a^{\prime})$. Moreover, $\eth a$ is an immersed $C^{0}$ submanifold of codimension 2. However, $\eth a$ need not be embedded (for example, in $2$ spatial dimensions, suppose that $a$ is the union of two disjoint wedges whose edges intersect at a point).
@@
The area of $\eth a$ will be denoted $\operatorname{Area}(a)$.
@@
Given a wedge $a$, we distinguish between its edge $\eth a$ in $M$ and its edge $\delta\hskip-1.99997pt\bar{}\hskip 1.99997pta$ as a subset of the conformal completion $\tilde{M}$ Wald:1984rg . If $a$ is asymptotic, the latter set may contain an additional piece, the conformal edge
$\tilde{\eth}a\equiv\delta\hskip-1.99997pt\bar{}\hskip 1.99997pta\cap\partial\tilde{M}~.$ (8)
@@ 3.1 Definitions
We emphasize that no numerical value is assigned to the expansion by the present definition; see Def. 40 below. To define past-nonexpanding, $H^{-}$ replaces $H^{+}$ in Eq. (21).
Definition 29 (Noncontraction).
A wedge $a$ is future-noncontracting if there exists an open set $O\supset\eth a$ such that no proper past-directed outward null deformation of $a$, with compact support within $O$, is future-nonexpanding on all new edge points. That is, no wedge $b\supsetneq a$ with $\eth b\subset\eth a\cup[H^{-}(a^{\prime})\cap O]$ is future-nonexpanding on all points in $\eth b\setminus\eth a$. (Heuristically, in sufficiently smooth settings, future-noncontracting implies nonnegativity of the outward future quantum expansion of $a$; moreover, positivity of the quantum expansion implies future-noncontracting.) Past-noncontracting is defined analogously.
@@
Let $\eth a^{+}$ ($\eth a^{-}$) denote the set of points $p\in\eth a$ where the wedge $a$ is future- (past-) nonexpanding.
@@
$\eth a^{+}$ and $\eth a^{-}$ are open subsets of $\eth a$ in the induced topology of $\eth a$.
@@
Let $p\in\eth a^{+}$. We must show that $\eth a^{+}$ contains an open subset that contains pp. Let $O$ be the open set guaranteed to exist by Def. 28, and let $q\in\eth a\cap O$. Then $O$ contains $q$ and satisfies Eq. (21). Hence all points in $\eth a\cap O$ are future-nonexpanding: $\eth a\cap O\subset\eth a^{+}$. ∎
@@
Let $a$ be a wedge and let $\eth a^{+}$ be the set of points where $a$ is future-nonexpanding. The future-lightsheet of $a$ is the null hypersurface
@@
Figure 2: Illustration of Definition 28: a wedge $a$ is future-nonexpanding at $p$ if there exists an open neighborhood $O$ of $p$ such that all future-outward deformations of $\eth a$ within $O$ (green) do not increase the area (classically), or (semiclassically) have nonpositive smooth max-generalized entropy conditioned on $a$.
Figure 3: Discrete Max-Focusing. If the wedge $a$ is future-nonexpanding on a portion $\eth a^{+}$ of its wedge $\eth a$, then $a$ can be outward deformed along the future lightsheet emanating from $\eth a^{+}$ into wedges $b$, $c$ that obey $H_{\rm max,gen}(c|b)\leq 0~.$
@@
Let $a$, $b$, and $c$ be wedges such that $a\subset b\subset c$, and suppose that $\eth b\cup\eth c\subset\eth a\cup L^{+}(a)$. Then
@@
The same statement holds in the past direction, i.e., if $\eth b\cup\eth c\subset\eth a\cup L^{-}(a)$.
@@ 3.2 Immediate Implications
By assumption there exists an open neighborhood $O$ of $p$ in which the past and future null vectors normal to $\eth a$ are tangent to $\eth a$. None of the null generators of $H^{\pm}(a^{\prime})$ start on this portion. Since $a$ cannot be deformed along $H^{\pm}(a^{\prime})$ at all, there exists no wedge $b$ for which Eq. (21) could fail. ∎
Figure 4: Applications of Definition 28. All wedges shown are understood to be subsets of Minkowski spacetime. Left: the wedge $a$ is nonexpanding at all the dark blue points, including at the (concave) corner (blue dot). Center: $a$ is future-nonexpanding except at the (convex) corner, where the green deformation $\eth a\to\eth b$ fails to satisfy Eq. (21). Right: the time direction is suppressed, but all three spatial directions are now displayed. The edge $\eth a$ is now shown as a shaded surface. The corner (red dot) is neither convex nor concave. Such a corner fails to be future-nonexpanding, because in any open neighborhood (orange), the green deformation increases the area.
@@
Let $a$ and $b$ be wedges such that $a\subset b$, and suppose that $\eth b\subset\eth a\cup L^{+}(a)$. Assuming Discrete Max-Focusing, Conjecture 33,
$\eth b\cap L^{+}(a)\subset\eth b^{+}~.$ (24)
The same statement holds in the past direction: ++ can be replaced by - throughout.
@@
By Lemma 31, $\eth a\setminus\eth a^{+}$ is closed in the induced topology of $\eth a$. Hence $\eth b\cap L^{+}(a)=\eth b\setminus(\eth a\setminus\eth a^{+})$ is open in the induced topology of $\eth b$. Therefore, there exists a set $O$ that is open in $M$ such that $\eth b\cap O=\eth b\cap L^{+}(a)$.
Let $c\supset b$ such that $\eth c\subset\eth b\cup[H^{+}(b^{\prime})\cap O]$. Since $H^{+}(b^{\prime})\subset H^{+}(a^{\prime})$, we have $\eth c\subset\eth a\cup L^{+}(a)$ and hence the assumptions of Discrete Max-Focusing, Conjecture 33, are satisfied. The Conjecture then implies Eq. (25). ∎
@@
Let $a\subset c$ such that $\eth c\subset\eth a\cup L^{+}(a)$. That is, $c$ is a future-outward deformation of $a$ along orthogonal lightrays originating on the future-nonexpanding portion of its edge $\eth a$. Then
@@
The same statement also holds if $\eth c\subset\eth a\cup L^{-}(a)$.
@@
Let $\cal C$} be the boundary of the past of a subset $s$ of the conformal boundary of an arbitrary spacetime $M$. (Note that $s$ can be, but need not be, a single point.) If there exists a Cauchy slice $\Sigma$ of $M$ such that any wedge $a$ with $\eth a\subset{\cal C}\cap J^{+}(\Sigma)$ is past-nonexpanding, then $\cal C$ is called a future-causal horizon.
@@
Let $\cal C$ be a future-causal horizon. Let $b\subset c$ be wedges whose edges both lie on $\cal C$ with $\eth c\subset L^{-}(b)$. Then